# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

### v7

@@ 3.5 Performance
Before assessing the utility of AGITB, it is essential to recognise its role as a pragmatic benchmark rather than an end in itself. Similar to the Turing Test—which Harnad1992 described as an empirical criterion rather than a philosophical sleight of hand—AGITB is intended as a practical instrument for evaluating progress toward artificial general intelligence. The ultimate objective remains the development of AGI, not merely success on the benchmark.
AGITB can yield meaningful insights only when developers adhere strictly to its core requirements. Misinterpretations of fundamental elements, such as the "empty" initial state, may lead to erroneous conclusions and hinder genuine progress toward AGI.
Overall, AGITB provides a structured framework for empirically assessing the AGI potential of diverse computational paradigms, including classical symbolic architectures, artificial neural networks, and large language models. However, before benchmarking artificial systems, it is necessary to consider the baseline performance of human cognition.
@@ 3.5.1 Human Performance
The inability to directly compare internal cortical states makes it impossible to verify AGITB requirements in humans in a strict computational sense. Nevertheless, because cortical architecture inherently supports low-level binary signal processing and the tests align with cognitive expectations, it can be assumed that humans naturally satisfy most requirements. Demand 1 (Unbiased start), however, warrants further discussion.
Owing to prior experience and cognitive bias, an adult human may appear to "fail" the unbiased start prerequisite, as the cortex is no longer in an unconditioned state and may generate non-empty predictions. However, AGITB specifies that the unbiased state must precede the first input, a condition met in the fetal cortex. At that stage, the human cortex lacks synaptic organisation and, upon receiving its first sensory input, would satisfy the criterion of true neutrality by predicting no spikes.
Several of the more complex AGITB tests have cognitive-level analogues that can be observed through conscious reasoning and introspection. Temporal flexibility (Requirement 7), for example, is straightforward for humans, who can easily recognise temporal patterns of varying durations. Stagnation (Requirement 8) reflects the physical limitation of the human cortex to store and maintain knowledge. The corresponding test is analogous to the onset of cognitive saturation or early dementia, when recent experiences are forgotten. Because humans learn different topics at varying speeds, the rationale for content sensitivity (Requirement 9) and context sensitivity (Requirement 10) is self-evident. The unobservability test (Requirement 11) captures the fact that internal human mental states cannot be directly inspected. The denoising test (Requirement 12) and the generalisation test (Requirement 13) correspond to cognitive abilities in which humans excel—recalling and generalising when faced with new or distorted inputs. Finally, bounded prediction latency (Requirement 14) is a defining property of the human brain.
@@ 3.5.2 Classical Symbolic Programs
In principle, two complementary designs for AGI can be conceived. The first encodes, either explicitly or implicitly, certain expectations about the external world within the system itself. The second is entirely expectation-free, beginning from a neutral state without built-in assumptions or prior knowledge.
The former category encompasses most AI and putative AGI systems developed to date, yet it remains fundamentally constrained by the Symbol Grounding Problem (SGP), as articulated by Harnad1990. Although such systems may exhibit behaviour that appears intelligent, they lack grounded understanding and therefore cannot be regarded as genuine instances of AGI.
More specifically, systems that embed expectations inevitably reflect the programmer’s assumptions about the meaning of the signals from which they learn. In the case of standard symbolic programming, the program itself constitutes prior knowledge—its rules and representations presuppose an interpretation of the symbols being manipulated. The existence of such a program directly violates AGITB’s first test, which prohibits the inclusion of any external knowledge. Namely, the AGI program itself introduces the symbol-grounding problem it is meant to overcome.
AGITB’s initial test thus formalises the requirement to learn from scratch, demanding that a system derive structure and meaning solely through exposure to intrinsically grounded binary inputs. Only an expectation-free system could, in principle, satisfy this condition. To date, however, no such program is known. This suggests that a true AGI may not explicitly encode the operations of intelligence itself, but rather the dynamics of the substrate from which intelligence can emerge—a view consistent with the "Brain Simulator Reply" proposed by Churchland1990 in response to Searle’s Chinese Room Argument (Searle1980).
@@ 3.5.3 Artificial Neural Network Performance
A similar dichotomy applies to connectionist architectures. Artificial neural networks can, in principle, be constructed either with built-in expectations—through pretraining and architectural priors—or as expectation-free systems that begin from a neutral state. However, AGITB’s requirement that a model start in a completely unbiased, knowledge-free condition stands in fundamental tension with the prevailing paradigm of modern deep learning. Contemporary neural models typically rely on pretraining, whereby network weights are adjusted in advance through exposure to structured or labelled data.
Only an untrained network can be regarded as unbiased. However, an untrained network with all weights initialised to zero exhibits no learning dynamics: the neurons are not halted but remain dormant, producing no output. This exposes a fundamental limitation of current ANN architectures—they lack an intrinsic mechanism for initiating unbiased, autonomous learning from scratch in arbitrary and unfamiliar environments.
Furthermore, by translating symbols into numerical vectors, ANNs effectively transform the symbol grounding problem into a number grounding problem. While these numerical representations capture relational regularities among symbols, they also introduce spurious associations that are not anchored in real-world semantics, giving rise to the phenomenon commonly referred to as hallucination.
Pretraining on large datasets establishes correlations rather than genuine understanding. AGITB rejects this approach in favour of internal learning algorithms capable of initiating adaptation autonomously. However, no such mechanism has yet been demonstrated.
@@ 3.5.4 Large Language Model Performance
Because a large language model (LLM) is a type of artificial neural network (ANN), it fails the unbiased start test (Requirement 1). This failure demonstrates that LLMs cannot learn entirely from scratch and are inherently biased toward the language and data on which they were trained. Although an LLM’s internal state is readily accessible—and could, unlike that of a human, be inspected for further testing—such evaluations remain moot until the first requirement is satisfied.
Two possibilities warrant consideration. The first concerns whether LLMs, while not qualifying as AGI, might nevertheless simulate one. For example, an LLM could in principle be prompted to imitate a model receiving binary inputs as specified by AGITB; a structured prompt of this type is provided in the Appendix. However, before evaluating such simulations, it is necessary to revisit the sensitivity requirement (Requirement 4), which stipulates that a model must remain responsive to any individual input even after processing an arbitrarily long sequence of subsequent inputs. This condition is fundamentally incompatible with the architecture of LLMs, which operate within a fixed context window. Once this window is exceeded, earlier tokens are discarded or compressed, causing the system to lose or attenuate the long-term dependencies required by the test.
The second possibility is whether an LLM can autonomously generate an architecture capable of satisfying AGITB, given the reference implementation. In practice, this has not been observed: although systems such as ChatGPT-5 produced candidate programs purporting to pass the tests, none advanced beyond the adaptability test (Requirement 7).
To summarise, LLMs do not perform genuine learning solely from prompts, nor can they develop the grounded, context-dependent understanding of the world characteristic of human cognition. These limitations extend to large reasoning models (LRMs), which share the same underlying architectural principles.
@@ 3.6 Remarks
AGITB evaluates the model’s predictive capabilities following exposure to temporal sequences of either structured or random inputs. Random input sequences with arbitrary internal correlations are used to minimise reliance on pretraining, ensuring that any observed learning arises from the input stream rather than prior knowledge. By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals—the random inputs need not resemble real-world sensory data.
The low-level, binary operational framework makes AGITB particularly well-suited for evaluating NeuroAI models that aim to satisfy the principles of the embodied Turing Test, as proposed, for example, by Zador2023, where cognitive understanding emerges from the integration of continuous sensory streams. The progression from raw signal prediction to high-level abstraction reflects the broader evolution of AI itself—from early perceptrons to large-scale models such as GPT.
Hand-engineered, task-specific systems have historically struggled to scale toward true generalisation, reinforcing the need for adaptive, learning-based architectures such as artificial neural networks. However, state-of-the-art AI networks rely heavily on pretraining with structured or symbolic data—a process that inevitably introduces bias and circumvents the grounding of meaning in raw sensory input. This raises a critical and unresolved question: how can such models begin learning directly from unstructured input, as required to satisfy the AGITB?
@@ 3.7 Cheating the Benchmark
Since AGITB tests are simple to solve in isolation, a plausible strategy for circumventing the benchmark would be to develop task-specific solutions and make the model deploy them selectively based on the task at hand.
While it is theoretically possible to infer the model’s assigned task and role by monitoring the number of instantiated model objects and their invoked methods, such an effort would serve only to subvert the benchmark rather than to advance AGI research. Although AGITB could incorporate additional obfuscation strategies—such as shuffling tests and redesigning the programming interface—these measures would not only deter cheating but also make tests less intelligible to human developers. This would undermine the testbed’s core objectives: transparency, readability, and interpretability.
For this reason, the AGITB reference implementation remains deliberately readable and accessible. To date, no non-AGI system has demonstrated the level of performance that AGITB demands. Unless a credible attempt to circumvent the benchmark emerges, there is no justification for introducing an obfuscated version of the test.
@@ 4 Competing Benchmarks
Among existing benchmarks, the Abstraction and Reasoning Corpus (ARC) introduced by Chollet2019 is most closely aligned in spirit, as it likewise emphasises generalisation over task-specific optimisation. A related effort is NeuroBench, which is designed to support the development of neuromorphic approaches to AGI. Both ARC and NeuroBench employ a variety of correctness and complexity metrics to differentiate among non-AGI systems, as they are primarily intended to distinguish weaker from stronger narrow models. In contrast, AGITB is designed to identify only true positive instances of AGI.
@@ 4.1 ARC
ARC presents visual reasoning tasks that require a model to infer novel transformations from a small set of input-output examples—such as recolouring, rearranging, or modifying spatial patterns in a spatial grid.
However, ARC implicitly assumes the presence of high-level cognitive priors, such as object permanence, spatial reasoning, numerical abstraction, and causal inference. These priors are left undefined, placing an ambiguous burden on the model architecture. In contrast, AGITB adopts a fundamentally different perspective: it treats the model under evaluation as a blank slate that must develop structure and function exclusively through interaction with temporally structured input. Low-level reflexes are not regarded as prerequisites for intelligence within this framework, but rather as evolutionary features of the subcortical "old brain," as described by (Hawkins2004, p. 66).
While ARC presumes temporal reasoning, it does not explicitly test it, as it uses only two images to demonstrate a transformation. In contrast, AGITB evaluates cognition as a dynamic, unfolding process over time. A model can only develop knowledge and predictive capabilities through time, not from disconnected before-and-after snapshots that lack the temporal continuity needed to infer causal structure. For instance, to learn about objects and to recognise the object transformation, such as "move left," a model under AGITB must observe multiple intermediate snapshots of the environment across time; the final image alone provides insufficient information to infer anything. Learning requires exposure to temporally structured data, enabling the model to discover invariants and causal relations through experience.
In summary, ARC requires a high-level intelligence, whereas AGITB requires the 14 low-level principles. ARC evaluates performance based on behaviours grounded in human-like cognitive priors, while AGITB determines a system’s ability to support the emergence of these priors through exposure to raw input. Benchmarks such as ARC and the Turing test implicitly frame intelligence through the lens of human cognition, thereby embedding anthropocentric biases into their evaluation criteria. In contrast, AGITB conceptualises intelligence as a universal capacity for learning independent of innate symbolic structures or species-specific assumptions.
@@ 4.2 NeuroBench
NeuroBench provides a unified framework for benchmarking diverse AI models across a standardised set of tasks and metrics. It is particularly specialised for neuromorphic approaches, which have demonstrated advantages in resource efficiency and scalability. Within its algorithm track, the framework evaluates model performance on several ongoing challenges in general AI research, including few-shot continual learning, object detection, sensorimotor decoding, and predictive modelling.
The predictive modelling challenge, which involves forecasting chaotic functions, is most closely aligned with AGITB’s central premise that intelligence fundamentally concerns the prediction of future states. For this purpose, NeuroBench employs a synthetic one-dimensional Mackey–Glass time series—a dataset specifically designed for architectures with limited input/output capacity.
However, several issues limit NeuroBench’s suitability as a general AGI benchmark. First, the Mackey–Glass data are numerical; the meaning of a number, much like that of a symbol, is defined by human interpretation. The use of such representations necessarily reintroduces the symbol-grounding problem. Second, the threshold corresponding to AGI-level performance remains undefined. Although the symmetric mean absolute percentage error (sMAPE) is an established metric in forecasting and NeuroBench provides a baseline score, it does not specify what level of accuracy constitutes general intelligence. After all, humans perform poorly at detecting anomalies or making long-term predictions of the Mackey–Glass signal (Thill2020). Third, although long-term prediction is not problematic in itself, predicting multiple steps ahead without timely feedback deprives a system of the opportunity to recognise and correct its own errors. This design is incompatible with online learning, where an AGI should continuously update its internal state upon detecting discrepancies between its predictions and the actual outcomes. NeuroBench, by contrast, implicitly assumes that an AGI would behave as a purely mechanical predictor, lacking intrinsic mechanisms for self-correction, autonomous adaptation, or genuine agency.

### v8

@@ 3.5 Performance
3.6 Performance
Before assessing the utility of AGITB, it is essential to recognise its role as a pragmatic benchmark rather than an end in itself. Similar to the Turing Test, which Harnad1992 characterised as an empirical criterion rather than a philosophical manoeuvre, AGITB is intended as a practical instrument for evaluating progress toward artificial general intelligence. The ultimate objective remains the development of AGI, not merely achieving success on the benchmark.
AGITB yields meaningful insights only when developers adhere strictly to its core requirements. Misinterpretations of fundamental elements, such as the notion of an “empty” initial state, may lead to erroneous conclusions and impede genuine progress toward AGI.
Overall, AGITB provides a structured framework for empirically assessing the AGI potential of diverse computational paradigms, including classical symbolic architectures, artificial neural networks, and large language models. Before benchmarking artificial systems, however, it is necessary to consider the baseline performance of human cognition.
@@ 3.6.1 Human Performance
The inability to directly compare internal cortical states makes it impossible to verify AGITB requirements in humans in a strict computational sense. Nevertheless, because cortical architecture inherently supports low-level binary signal processing and the tests align with basic cognitive competencies, it is reasonable to assume that humans naturally satisfy most requirements. Demand 1 (Unbiased start), however, warrants further discussion.
Owing to prior experience and cognitive bias, an adult human’s cortex may appear to “fail” the unbiased start prerequisite, as it is no longer in an unconditioned state and may generate non-empty predictions. AGITB, however, requires that the unbiased state precede the first input—a condition satisfied only in the fetal cortex. At that stage, the cortex lacks synaptic organisation and, upon receiving its first sensory input, meets the criterion of true neutrality.
Several of the more complex AGITB tests have cognitive-level analogues that can be observed through reasoning and introspection. Temporal flexibility (Requirement 7), for example, poses no difficulty for humans, who readily recognise temporal patterns of varying durations. Stagnation (Requirement 8) reflects the finite capacity of the human cortex to store and maintain knowledge; its behavioural analogue resembles the onset of cognitive saturation or early dementia, in which recent experiences are lost. Because humans learn different types of content at varying rates, the rationale for content sensitivity (Requirement 9) and context sensitivity (Requirement 10) is immediately evident. The unobservability test (Requirement 11) formalises the fact that internal human mental states cannot be directly inspected.
The denoising test (Requirement 12) and the generalisation test (Requirement 13) correspond to cognitive abilities in which humans excel—recalling and generalising when confronted with new or distorted inputs. Finally, bounded prediction latency (Requirement 14) is a well-established property of the human brain, which maintains approximately constant reaction times through extensive parallel processing.
@@ 3.6.2 Classical Symbolic Programs
In principle, two complementary designs for AGI can be conceived. One embeds, either explicitly or implicitly, expectations about the external world within the system. The other is entirely expectation-free, beginning from a neutral state without built-in assumptions or prior knowledge.
The former category includes most AI and purported AGI systems developed to date, yet it remains fundamentally constrained by the Symbol Grounding Problem (SGP), as articulated by Harnad1990. Although such systems may display behaviour that appears intelligent, their interpretations of symbols depend on programmer-supplied conventions rather than grounded understanding, and they therefore cannot qualify as genuine AGI.
More specifically, systems that embed expectations inevitably incorporate the designer’s assumptions about the meaning of the signals they process. In classical symbolic architectures, the program itself constitutes prior knowledge: its rules and representations presuppose interpretations of the symbols being manipulated. The very existence of such a program violates AGITB’s first test, which prohibits external knowledge of any kind. In effect, the AGI program smuggles in the symbol-grounding problem it is meant to avoid.
AGITB’s initial test thus formalises the requirement to learn from scratch, demanding that a system derive structure and meaning solely through exposure to intrinsically grounded binary signals. Only an expectation-free system could, in principle, satisfy this condition. This perspective suggests that a genuine AGI may not explicitly encode the operations of intelligence, but rather the dynamics of a substrate from which intelligence can emerge—a view aligned with the “Brain Simulator Reply” proposed by Churchland1990 in response to Searle’s Chinese Room Argument (Searle1980). To date, however, no such system is known.
@@ 3.6.3 Artificial Neural Network Performance
A similar dichotomy applies to connectionist architectures: ANNs can, in principle, be constructed either with built-in expectations (introduced through pretraining regimes or architectural priors) or as expectation-free systems that begin from a neutral initial state.
AGITB’s requirement that models begin in a completely unbiased, knowledge-free state stands in fundamental tension with the dominant paradigm of modern deep learning. Contemporary neural models typically rely on extensive pretraining, during which network weights are shaped by prior exposure to structured or labelled data. Moreover, by mapping symbolic inputs to numerical vectors, standard ANNs effectively shift the symbol-grounding problem into a number-grounding problem. Although these vector representations capture relational regularities within the training data, they also introduce spurious associations not anchored in real-world semantics, leading to the phenomenon commonly described as hallucination.
Only an expectation-free network can be regarded as unbiased. Yet a network without any initialised weights exhibits no effective learning dynamics: its units are not halted, but remain dormant and produce no informative output. This highlights a fundamental limitation of current ANN architectures: they do not initiate learning autonomously but instead depend on an external training procedure to drive adaptation. AGITB, by contrast, requires a blank system capable of initiating adaptation autonomously in an unfamiliar environment. To date, no such mechanism has been demonstrated in artificial neural networks.
@@ 3.6.4 Large Language Model Performance
Because a large language model is a type of artificial neural network, it fails the unbiased start test (Requirement 1). LLMs cannot learn entirely from scratch, as their behaviour is shaped in advance by the linguistic patterns and data on which they were pretrained. Although an LLM’s internal state is readily accessible and, unlike that of a human, can be inspected or compared across instances, any such evaluations remain moot until the first requirement is satisfied.
Two possibilities warrant consideration. The first is whether LLMs, while not constituting AGI, might nevertheless simulate one. For example, an LLM could be prompted to imitate a model receiving binary inputs as specified by AGITB. However, before examining such simulations, it is necessary to consider the sensitivity requirement (Requirement 4), which stipulates that a model must remain responsive to any individual input even after processing an arbitrarily long sequence of subsequent inputs. This requirement is fundamentally incompatible with the architecture of LLMs, which operate within a fixed context window. Once this window is exceeded, earlier tokens are discarded or compressed, weakening or eliminating the long-term dependencies required by the test.
The second possibility is whether an LLM can autonomously generate a solution to AGITB when provided with the reference implementation. In practice, this has not been observed: although systems such as ChatGPT-5 produced candidate programs purported to satisfy the requirements, none progressed beyond the adaptability test (Requirement 7).
In summary, LLMs do not perform genuine learning solely from prompts, nor can they acquire the grounded, context-dependent understanding characteristic of human cognition. These limitations extend to large reasoning models, which share the same underlying architectural foundations.
@@ 3.7 Remarks
AGITB evaluates a model’s predictive capabilities after exposure to temporal sequences of both structured and random inputs. Random input sequences with arbitrary internal correlations are employed to minimise reliance on pretraining, ensuring that any observed learning arises from the input stream itself rather than from prior knowledge. By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals; the random inputs need not resemble real-world sensory data.
The low-level, binary operational framework makes AGITB particularly well suited for evaluating NeuroAI models that aim to satisfy principles of the embodied Turing Test, as proposed by Zador2023, wherein cognitive understanding emerges from the integration of continuous sensory streams. The progression from raw signal prediction to higher-level abstraction mirrors the broader trajectory of AI, from early perceptrons to large-scale models such as GPT.
Hand-engineered, task-specific systems have historically struggled to scale toward genuine generalisation, underscoring the need for adaptive, learning-based architectures such as artificial neural networks. However, state-of-the-art neural systems rely heavily on pretraining with structured or symbolic data, a process that inevitably introduces biases and bypasses the grounding of meaning in raw sensory input. This raises a central and unresolved question: how can such models initiate learning directly from unstructured input, as required to satisfy AGITB?
@@ 3.8 Cheating the Benchmark
Because AGITB’s tests are individually simple to solve, one might imagine circumventing the benchmark by engineering task-specific solutions and having the model selectively deploy them depending on the detected test scenario. In principle, the task being administered could be inferred by monitoring the number of instantiated models and the sequence of invoked methods.
However, such an approach would amount to subverting the benchmark rather than advancing AGI research. Although AGITB could be hardened against this form of cheating (by, for example, shuffling tests or redesigning the programming interface) these measures would reduce the transparency and interpretability of the testbed and thereby hinder its intended use by human developers.
For this reason, the AGITB reference implementation is kept deliberately readable and accessible. To date, no artificial system has demonstrated the level of performance that AGITB requires. Unless a credible attempt to circumvent the benchmark emerges, there is no justification for introducing a more obfuscated version of the testbed.
@@ 4 Competing Benchmarks
Among existing benchmark tasks, the Abstraction and Reasoning Corpus (ARC) introduced by Chollet2019 is most closely aligned in spirit, as it likewise emphasises generalisation over task-specific optimisation. A related effort is NeuroBench, which is designed to support the systematic evaluation of neuromorphic and other biologically inspired architectures. Both ARC and NeuroBench rely on a variety of correctness and complexity metrics to compare non-AGI models; their primary purpose is to distinguish weaker from stronger narrow systems. In contrast, AGITB is constructed to assess whether a model satisfies a set of necessary conditions associated with artificial general intelligence, rather than to rank models within a specific performance spectrum.
@@ 4.1 ARC
ARC presents visual reasoning tasks in which a model must infer novel transformations (such as recolouring, rearranging, or modifying spatial patterns) from a sequence of two input–output examples defined on discrete spatial grids.
However, ARC implicitly assumes the presence of high-level cognitive priors, including object permanence, spatial reasoning, numerical abstraction, and causal inference. These priors are not formally specified, placing an ambiguous and open-ended burden on the model designer. In contrast, AGITB adopts a fundamentally different stance: it treats the system under evaluation as a blank slate that must acquire structure and function exclusively through interaction with temporally structured input.
Although ARC presumes temporal reasoning, it does not explicitly test it, since each task presents only two images to illustrate a transformation. AGITB, by contrast, evaluates cognition as a dynamic process unfolding over time. A model can acquire knowledge and predictive capability only through continuous exposure to temporally structured data, not from disconnected before–after snapshots that lack the temporal continuity needed to infer causal relationships. For example, to recognise an object moving left, a model in AGITB must observe multiple intermediate states across time; the final image alone is insufficient to infer the transformation. Temporal structure, rather than static pattern comparison, provides the substrate from which invariants and causal relations can be learned.
In summary, ARC evaluates high-level intelligence grounded in human cognitive priors, whereas AGITB evaluates adherence to fourteen low-level computational principles intended to support the emergence of such priors. ARC and the Turing Test both frame intelligence through an anthropocentric lens, embedding assumptions drawn from human cognition. AGITB instead conceptualises intelligence as a universal capacity for learning that does not rely on innate symbolic structures or species-specific expectations.
@@ 4.2 NeuroBench
NeuroBench provides a unified framework for benchmarking diverse AI models across a standardised set of tasks and metrics. It is particularly oriented toward neuromorphic approaches, which have demonstrated advantages in resource efficiency and scalability. Within its algorithm track, the framework evaluates models on several challenges relevant to general AI research, including few-shot continual learning, object detection, sensorimotor decoding, and predictive modelling.
The predictive modelling challenge, which involves forecasting chaotic functions, is most closely aligned with AGITB’s central premise that intelligence fundamentally concerns the prediction of future states. NeuroBench employs a synthetic one-dimensional Mackey-Glass time series for this task, a dataset designed for architectures with limited input/output capacity.
However, several issues limit the usability of chaotic function prediction (CFP) as a general AGI benchmark task. First, the Mackey-Glass data are numerical, and NeuroBench does not prescribe the encoding scheme. An inappropriate encoding can distort the temporal and causal structure of the observed signals, such that a numeric value—much like a symbol—derives its meaning from human interpretation rather than from the model’s own grounded understanding. This effectively reintroduces the symbol-grounding problem in a numerical form.
Second, the threshold for AGI-level performance is not clearly defined. Although symmetric mean absolute percentage error (sMAPE) is a standard forecasting metric, NeuroBench does not specify what performance level would correspond to general intelligence. Notably, humans themselves perform poorly at anomaly detection and long-horizon prediction of the Mackey-Glass signal (Thill2020).
Third, although long-term prediction is not inherently problematic, predicting multiple steps ahead without timely feedback deprives a system of the opportunity to detect and correct its own errors. This design is incompatible with online learning, in which an AGI should continuously update its internal state upon observing discrepancies between prediction and outcome. NeuroBench, by contrast, emphasises offline learning and assumes that an AGI would behave as a purely mechanistic predictor, lacking intrinsic mechanisms for self-correction, autonomous adaptation, and genuine agency.